The current price of a stock is $ 53.15 and the annual risk-free rate is 6.6 percent. A put option with an exercise price of $55 and one year until expiration has a current value of $ 4.98 . What is the value of a call option written on the stock with the same exercise price and expiration date as the put option?

Note, the given interest rate is an effective rate, so for calculation purposes, you need only discount the using the risk free rate, no e x adjustment is needed.
Answer #1

Using Put Call parity,

C + X/(1 + r)n = S + P

C + 55/(1.066) = 53.15 + 4.98

C = $6.54

Price of Call Option = $6.54